# The number of spanning trees as an indicator of critical phenomena: When Kirchhoff meets Ising

### v1

@@ Abstract
I Introduction
Capturing physical properties from graphs or networks has been extensively explored in Statistical Mechanics [1, 2]. Among the many concepts related to these two terms (here, the only difference between them lies in the level of abstraction and the context of application), the concept of a tree is particularly important (for an excellent textbook on graph theory, see Ref. [3]). A tree is the simplest way of connecting all nodes in a graph without redundancies, that is, without cycles. In other words, a tree is a minimal structure in the sense that removing any edge would break the connectivity of the graph.
@@ I Introduction
This framework enables the topological quantification of temporal complexity and the extraction of scaling properties directly from the network structure. But what about time series generated by spin system models? Zhao et al. [19] investigated networks constructed from time series of the Ising model, yet without exploring the spectral properties of random matrices represented by adjacency matrices or, more importantly, addressing the main question of the present work: what can the number of spanning trees of graphs constructed from different evolutions of spin systems governed by the traditional Monte Carlo Markov Chain (MCMC) prescribed by the Metropolis algorithm reveal about the phase transitions of such systems?
@@
II Spectral signatures of spin dynamics through visibility graphs
Let us define
@@ II Spectral signatures of spin dynamics through visibility graphs
Several studies have shown that, during the early stages of evolution, criticality can be identified through temporal power laws [21, 22, 23], from which all critical exponents of different models can be determined. For example, for $m_{0j}=1$, it is expected that, at $T=T_{C}$ and for sufficiently large systems, $m_{i}=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{ij}\sim i^{-\frac{\beta}{\nu z}}$, where $\beta$ and $\nu$ are the standard static exponents, and $z$ is the dynamic exponent. For $T\neq T_{C}$, deviations from this power law are typically observed and are well described by a stretched exponential.
On the other hand, for simulations starting with a small but fixed initial magnetization, $m_{0j}=\varepsilon\ll 1$, the system exhibits a power-law behavior characterized by a new dynamic exponent $\theta$, such that $m_{i}\sim i^{\theta}$. This exponent can also be determined without fixing the initial magnetization–i.e., by considering random initial configurations ($T\rightarrow\infty$)–through averaging over the initial magnetizations using the total correlation $c_{i}=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{ij}m_{0j}$, where $\langle m_{0j}\rangle=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{0j}\approx 0$. In this case, the same power-law behavior is observed, namely $c_{i}\sim i^{\theta}$ [24].
@@
The eigenvalues of symmetric matrices of dimension $N$, whose entries are random variables identically distributed according to a symmetric probability density function (PDF) with all moments finite, necessarily follow Wigner’s semicircle law: $\rho(\lambda)=\frac{\mathbf{1}_{\left|\lambda\right|\leq 2\sigma\sqrt{N}}}{2\pi N\sigma^{2}}\sqrt{4N\sigma^{2}-\lambda^{2}}$, where $\sigma^{2}$ is the second moment of the matrix entries and $\mathbf{1}_{\left|\lambda\right|\leq R}$ is the indicator function, which takes the value 1 for $\left|\lambda\right|\leq R$ and 0 otherwise.
Although adjacency matrices $A$ of random graphs, when standardized as $A_{s}=\frac{A-pJ}{\sqrt{Np(1-p)}}$, where $J_{ij}=1$ for $i,j=1,\ldots,N$, have eigenvalues whose bulk converges (as $N\rightarrow\infty$) to Wigner’s semicircle law, $\rho(\lambda)=\frac{\mathbf{1}_{\left|\lambda\right|\leq 2}}{2\pi}\sqrt{4-\lambda^{2}}$, in the dense regime (where $p$ is constant) one eigenvalue escapes from the bulk with a pronounced value $\lambda(A)=Np+o(N)$. For the intermediate case, when $Np$ is of the order of a few tens, the same centering and normalization yield a good fit to the semicircle law, although small fluctuations may occur. Finally, in the sparse regime, $p=c/N$ with $c$ constant, the graph becomes sparse and behaves similarly to a random tree, since its construction follows a Galton–Watson stochastic (branching) process. In this case, the histogram of eigenvalues no longer follows the semicircle law: there is a concentration – or aggregation – near zero and a flattened bulk, accompanied by the occurrence of outliers outside this interval.
@@
Kirchhoff’s Theorem (Matrix–Tree Theorem): All cofactors of the Laplacian matrix $L$ are equal and correspond to the number of spanning trees of $G$.
@@
III Results
We begin our results section by presenting examples of visibility graphs constructed from time series of the magnetization per spin of the Ising model, evolved via MCMC simulations using the Metropolis algorithm. The time evolutions of the system for two different temperatures $T=0.45$ and $T=15.6$, each for three distinct random seeds, are shown in Fig. 1.
@@ III Results
At low temperature ($T=0.45$), below the critical one $T_{C}=(1/2)ln(1+sqrt(2))$ the strong spin-spin correlations in the ordered Ising system give rise to slow, collective fluctuations of the magnetization. The corresponding visibility graphs display hubs, higher clustering, and a more heterogeneous degree distribution, reflecting the strong temporal correlations. In contrast, at high temperature ($T=15.6$), the magnetization fluctuates around zero with almost no temporal correlation. The visibility graphs are more homogeneous, with a narrow degree distribution and low clustering, resembling a random network, which captures the increased stochasticity of the dynamics.
@@
Shorter tails are observed at higher temperatures and longer ones at lower temperatures, consistent with the behavior of the eigenvalue density. The case $T=T_{C}$ is highlighted for reference. Qualitatively, the eigenvalue density of visibility graphs constructed from the temporal evolution of magnetization appears to reflect the system’s thermodynamics. Similar spectral signatures have been reported in the analysis of correlation matrices (see, e.g., [28, 29, 30, 31]). We next examine whether the logarithm of the number of spanning trees, defining the structural entropy, can serve as a thermodynamic indicator. In particular, we test whether the number of spanning trees responds to the model’s phase transition.
@@
IV Conclusions
Our spectral analysis of visibility graphs constructed from the time evolution of magnetization shows that the number of spanning trees, or simply its logarithm, provides a nonconventional indicator of the thermodynamics of spin systems and captures the phase transition of the Ising model, similarly to the behavior observed in the histograms of eigenvalues and eigenvalue spacings of the adjacency matrices. It is, of course, necessary to reinterpret the meaning of this “spectral thermodynamics,” particularly regarding the connection between criticality and the maximum of the structural entropy (also referred to as tree entropy when divided by the number of nodes in the thermodynamic limit; see, for example, [39]).
@@ IV Conclusions
Acknowledgments
RDS, HAF, and AA thank CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) of Brazil for financial support under grants 304575/2022-4, 405508/2021-2, and 307317/2021-8, respectively. RDS also thanks Prof. M. J. de Oliveira (IF-USP) for his valuable insights.

### v2

@@ Abstract
INTRODUCTION: Capturing physical properties from graphs or networks has been extensively explored in Statistical Mechanics [1, 2]. Among the many concepts related to these two terms (here, the only difference between them lies in the level of abstraction and the context of application), the concept of a tree is particularly important (for an excellent textbook on graph theory, see Ref. [3]). A tree is the simplest way of connecting all nodes in a graph without redundancies, that is, without cycles. In other words, a tree is a minimal structure in the sense that removing any edge would break the connectivity of the graph.
@@
This framework enables the topological quantification of temporal complexity and the extraction of scaling properties directly from the network structure. But what about time series generated by spin system models? Zhao et al. [19] investigated networks constructed from time series of the Ising model, and Moraes and Ferreira [20] similar results on the contact process, both yet without exploring the spectral properties of random matrices represented by adjacency matrices or, more importantly, addressing the main question of the present work: what can the number of spanning trees of graphs constructed from different evolutions of spin systems governed by the traditional Monte Carlo Markov Chain (MCMC) prescribed by the Metropolis algorithm reveal about the phase transitions of such systems?
@@
SPECTRAL SIGNATURES OF SPIN DYNAMICS THROUGH VISIBILITY GRAPHS: Let us define
@@
Several studies have shown that, during the early stages of evolution, criticality can be identified through temporal power laws [22, 23, 24, 25, 26], from which all critical exponents of different models can be precisely determined [24, 25, 27, 28]. For example, for $m_{0j}=1$, it is expected that, at $T=T_{C}$ and for sufficiently large systems, $m_{i}=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{ij}\sim i^{-\frac{\beta}{\nu z}}$, where $\beta$ and $\nu$ are the standard static exponents, and $z$ is the dynamic exponent. For $T\neq T_{C}$, deviations from this power law are typically observed and are well described by a stretched exponential.
On the other hand, for simulations starting from a small but fixed initial magnetization, $m_{0j}=\varepsilon\ll 1$, the system exhibits a power-law behavior characterized by a new dynamic exponent $\theta$, such that $m_{i}\sim i^{\theta}$ (see, for example, Refs. [29, 30, 31, 28]).
This exponent can also be determined without fixing the initial magnetization–i.e., by considering random initial configurations ($T\rightarrow\infty$)–through averaging over the initial magnetizations using the total correlation $c_{i}=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{ij}m_{0j}$, where $\langle m_{0j}\rangle=\frac{1}{N_{run}}\sum_{j=1}^{N_{run}}m_{0j}\approx 0$. In this case, the same power-law behavior is observed, namely $c_{i}\sim i^{\theta}$ [32].
@@
The eigenvalues of symmetric matrices of dimension $N$, whose entries are random variables identically distributed according to a symmetric probability density function (PDF) with all moments finite, necessarily follow Wigner’s semicircle law [33]: $\rho(\lambda)=\frac{\mathbf{1}_{\left|\lambda\right|\leq 2\sigma\sqrt{N}}}{2\pi N\sigma^{2}}\sqrt{4N\sigma^{2}-\lambda^{2}}$, where $\sigma^{2}$ is the second moment of the matrix entries and $\mathbf{1}_{\left|\lambda\right|\leq R}$ is the indicator function, which takes the value 1 for $\left|\lambda\right|\leq R$ and 0 otherwise.
Although adjacency matrices $A$ of random graphs, when standardized as $A_{s}=\frac{A-pJ}{\sqrt{Np(1-p)}}$, where $J_{ij}=1$ for $i,j=1,\ldots,N$, have eigenvalues whose bulk converges (as $N\rightarrow\infty$) to Wigner’s semicircle law, $\rho(\lambda)=\frac{\mathbf{1}_{\left|\lambda\right|\leq 2}}{2\pi}\sqrt{4-\lambda^{2}}$, in the dense regime (where $p$ is constant) one eigenvalue escapes from the bulk with a pronounced value $\lambda(A)=Np+o(N)$ [34, 35]. For the intermediate case, when $Np$ is of the order of a few tens, the same centering and normalization yield a good fit to the semicircle law, although small fluctuations may occur [36]. Finally, in the sparse regime, $p=c/N$ with $c$ constant, the graph becomes sparse and behaves similarly to a random tree, since its construction follows a Galton–Watson stochastic (branching) process. In this case, the histogram of eigenvalues no longer follows the semicircle law: there is a concentration – or aggregation – near zero and a flattened bulk, accompanied by the occurrence of outliers outside this interval [37, 38, 1].
@@
Kirchhoff’s Theorem (Matrix–Tree Theorem, see for example [3]). All cofactors of the Laplacian matrix $L$ are equal, and each of them corresponds to the number of spanning trees of the graph $G$.
@@
RESULTS: We begin our results section by presenting examples of visibility graphs constructed from time series of the magnetization per spin of the Ising model, evolved via MCMC simulations using the Metropolis algorithm. The time evolutions of the system for two different temperatures $T=0.45$ and $T=15.6$, each for three distinct random seeds, are shown in Fig. 1.
@@
At low temperature ($T=0.45$), below the critical one $T_{C}=\tfrac{1}{2}\ln(1+\sqrt{2})$ the strong spin-spin correlations in the ordered Ising system give rise to slow, collective fluctuations of the magnetization. The corresponding visibility graphs display hubs, higher clustering, and a more heterogeneous degree distribution, reflecting the strong temporal correlations. In contrast, at high temperature ($T=15.6$), the magnetization fluctuates around zero with almost no temporal correlation. The visibility graphs are more homogeneous, with a narrow degree distribution and low clustering, resembling a random network, which captures the increased stochasticity of the dynamics.
@@
Shorter tails are observed at higher temperatures and longer ones at lower temperatures, consistent with the behavior of the eigenvalue density. The case $T=T_{C}$ is highlighted for reference. Qualitatively, the eigenvalue density of visibility graphs constructed from the temporal evolution of magnetization appears to reflect the system’s thermodynamics. Similar spectral signatures have been reported in the analysis of correlation matrices (see, e.g., [40, 41, 42, 43, 44, 45]). We next examine whether the logarithm of the number of spanning trees, defining the structural entropy, can serve as a thermodynamic indicator. In particular, we test whether the number of spanning trees responds to the model’s phase transition.
@@
CONCLUSIONS: To summarize, we propose a novel analysis of visibility graphs constructed from the time evolution of magnetization shows that the number of spanning trees, or simply its logarithm, provides a nonconventional indicator of the thermodynamics of spin systems and captures the phase transition of the Ising model, similarly to the behavior observed in the histograms of eigenvalues and eigenvalue spacings of the adjacency matrices —a perspective that, to our knowledge, has not been explored before. It is, of course, necessary to reinterpret the meaning of this “spectral thermodynamics,” particularly regarding the connection between criticality and the maximum of the structural entropy (also referred to as tree entropy when divided by the number of nodes in the thermodynamic limit; see, for example, [51]).
@@
ACKNOWLEDGEMENTS: RDS, HAF, SG, and AA thank CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) of Brazil for financial support under grants 304575/2022-4, 405508/2021-2, 309560/2025-0, and 307317/2021-8, respectively. RDS also thanks Prof. M. J. de Oliveira (IF-USP) for his valuable insights.